OpenStudy (anonymous):

Trying to simplify this equation: sqrt 125z^8 / sqrt 16z^4

OpenStudy (anonymous):

\[\frac{ \sqrt{125}z^8 }{ \sqrt{16}z^4 } \]

OpenStudy (anonymous):

The above equation is the proper format if it's easier to see this way. Been stuck trying to simplify.

OpenStudy (mathstudent55):

Equation implies an equal sign. 2x + 5 = 15 is an equation. You have an expression that needs to be simplified.

OpenStudy (mathstudent55):

Do you know the rules of dividing powers?

OpenStudy (mathstudent55):

\(\large \dfrac{a^m}{a^n} = a^{m - n} \)

OpenStudy (anonymous):

I'm familiar with it, but idk how it applies towards radical equations. I looked into the answer key, and it says the answer is \[\frac{ 5z^2\sqrt{5} }{ 4 }\]

OpenStudy (mathstudent55):

That cannot possibly be the answer to the problem you posted. Is there an error in the problem? Are the z's supposed to be inside the roots?

OpenStudy (mathstudent55):

Is this th eproblem? \(\Large \frac{ \sqrt{125z^8} }{ \sqrt{16z^4} }\)

OpenStudy (anonymous):

correct